OpenStudy (anonymous):

find the equation of the line passing through the point (2, -4) and parallel to y = 3x - 6.

OpenStudy (amistre64):

strip the constant and recalibrate with the given point

OpenStudy (anonymous):

what?

OpenStudy (lalaly):

parallel means it has the same slope

OpenStudy (lalaly):

\[y-(-4)=3(x-2)\]\[y+4=3x-6\]\[y=3x-10\]

OpenStudy (amistre64):

y = 3x - 6 ; strip the constant y = 3x + C ; recalibrate with the given point -4 = 3(2) + C ; solve for C -10 = C ; rewrite y = 3x -10 yay!!

OpenStudy (amistre64):

parallel lines have the same equations but differ by a constant, that is if they are seperate lines

OpenStudy (lalaly):

yay!! xD